OpenStudy (anonymous):

Solve usingelimination method. If the ststem has not solution or an infinite number of solutions, state this -3x-8y=-88 and -3x-2y=-40

OpenStudy (anonymous):

can someone please help me with this one

OpenStudy (anonymous):

Solve using elimination method. If the system has no solution or an infinite number of solutions, state this: -3x-8y=-88 -3x-2y=-40 Will someone help me with this one. is this right -6x+10y=-48 -6x+10y-6=-48-6 10y=-48+-6 10y/10=-54/10 y=-54/10 then 6x-10y=-48 6x-10y-10=-48-10 6x=-48+-10 6x/6=-58/6 x=-58/6 will someone please tell me if this is right

OpenStudy (anonymous):

infinite solutions, right?

OpenStudy (kropot72):

Subtracting the second given equation from the first to eliminate x gives: -6y = -48 y = 6 Sustituting the value just obtained for y in the second given equation we get: 3x -(2 * 8) = -40 3x - 16 = -40 3x = -24 x = -8

OpenStudy (anonymous):

The solutions are x=y =8 You must have had some mis calculations above.

OpenStudy (anonymous):

-3x-8y=-88 -3x-2y=-40 -3x-8y=-88 +3x+2y=40 -6y = - 48 y =8 replace to find x=8

OpenStudy (kropot72):

Sorry. In the calculation of x I should have put -3x at the beginning and continued with the negative sign. The value of x is found from: -3x = -24 x = 8

OpenStudy (anonymous):

x=y=8

OpenStudy (anonymous):

ok so i am confused, can you rewrite the entire problem

OpenStudy (anonymous):

-3x-8y=-88 gives -3x=8y-88 put it in another equation ie -3x-2y=-40 we get y=8 and from this we get -3x=8(8)-88 -3x=-24 x=8

OpenStudy (anonymous):

and so that it the answer= x=8 what about y

OpenStudy (anonymous):

explain how you got this im confused

OpenStudy (phi):

If you have time, this video and the ones after it walk you through how to do this http://www.khanacademy.org/math/algebra/systems-of-eq-and-ineq/v/addition-elimination-method-1

OpenStudy (phi):

But, to answer your question, if you manage to find x=8, you would find y by using one of the equations, say this one -3x-8y=-88 replace x with 8: -3*8-8y= -88 simplify: -24-8y= -88 add 8y to both sides -24= 8y-88 add +88 to both sides 88-24= 8y 64= 8y divide both sides by 8 8=y

OpenStudy (anonymous):

k can you help me solve. Solve using the substitution method. show your work. If the system has no solution or an infinite number of solutions, state this. -7x+7y=-7 and x+9y=61

OpenStudy (phi):

the substitution method that means first solve for one of the variables. Pick x+9y=61 can you solve for x (by "moving" 9y to the other side)?

OpenStudy (anonymous):

i dont know can you

OpenStudy (phi):

can you add -9y to both sides?

OpenStudy (anonymous):

if you just will work this problem in full. i can figure it out from there. I am a person that learns by seeing

OpenStudy (anonymous):

just narrate as you work the problem. thanks

OpenStudy (phi):

x+9y=61 add -9y to both sides

OpenStudy (anonymous):

is it x=52

OpenStudy (phi):

you must not like letters. you have to keep -9y (can not combine with a number) x= 61-9y the y is important.

OpenStudy (anonymous):

the whole problem is -7x+7y=-7 and x+9y=61

OpenStudy (anonymous):

so 52y

OpenStudy (anonymous):

please just start from the beginning and writeout the whole solution to this problem. I am going to go to the site you gave me and evaluate what you gave me to better understand the whole thing

OpenStudy (phi):

you have to remember a rule: if you have 2 numbers you can combine them example 1+2=3 if you have a letter and a number example 61 -9y you can not combine them.

OpenStudy (phi):

so all we can do with x+9y=61 is this x+9y-9y= 61-9y we can simplify 9y-9y (we have 9 y's and subtract away 9 y's) = 0 so x= 61-9y Your 2 equations are -7x+7y=-7 x+9y=61 <-- this one we re-wrote as x= 61-9y now let's look at the first equation. Can you divide -7x + 7y = -7 by 7? (divide both sides (all terms)) by 7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

no i mean =2

OpenStudy (phi):

-7x + 7y = -7 divide each term by 7

OpenStudy (anonymous):

do i divide each number or add 7x and 7y together first to get 14 then divide by 7 then -7/7

OpenStudy (anonymous):

im sorry i am confused it would help me more to write the whole problem out and solution and solve and instruct each step all at once.

OpenStudy (phi):

You divide each term by 7. pay attention! 7x+7y IS NOT 14 7+7=14 I'll do the first term: -7x / 7 = -x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im sorry

OpenStudy (phi):

-7x + 7y = -7 divide each term by 7

OpenStudy (anonymous):

-x=-7 ? im sorry its been a long time for me

OpenStudy (phi):

we want to divide each term by 7 -7x/7 + 7y/7 = -7/7 (If we do the same thing to each term, the equation remains true) we get -x + y = -1 now we use the 2nd equation (remember it?) x= 61-9y this equation says x is the same as (61-9y) so replace the x in -x + y = -1 with (61-9y)

OpenStudy (anonymous):

is that it

OpenStudy (phi):

the next step is replace x in -x + y = -1 with (61-9y) can you write down what we get?

OpenStudy (anonymous):

hold on sorry had to go to rr

OpenStudy (anonymous):

52y=-1?

OpenStudy (phi):

don't do anything fancy. Just replace x with (61-9y) in -x + y = -1 and post it

OpenStudy (anonymous):

(61-9y)+y=-1

OpenStudy (phi):

what happen to the minus sign in front?

OpenStudy (anonymous):

52y+y=-1

OpenStudy (anonymous):

(61-(-9y)+y=-1

OpenStudy (phi):

you start with -x + y = -1 now replace x with (61-9y) -(61-9y) +y = -1 Do you see how to do that?

OpenStudy (anonymous):

you got your work cut out with me, Huh? lol

OpenStudy (anonymous):

why did you put the negative sign before the parenthesis

OpenStudy (anonymous):

oh oh im sorry i see now

OpenStudy (phi):

Now the minus sign is short hand for -1 times (stuff) . It is the same as -1*(61-9y) + y = -1 The next step is "distribute the -1" That means multiply every term inside the parentheses by -1. Can you do that?

OpenStudy (anonymous):

-61 and -9

OpenStudy (phi):

-1*(61-9y) = -1*61 + -1*-9y can you simplify (and stop ignoring the y)

OpenStudy (anonymous):

-52y=-60-9y= -52y=-51

OpenStudy (anonymous):

y sorry

OpenStudy (anonymous):

-52y=-51y

OpenStudy (phi):

for practice, simplify this -1*61 + -1*-9y

OpenStudy (anonymous):

-61+-9y=-70

OpenStudy (phi):

-61 is good but -1*-9y is +9y ( minus times a minus becomes a plus) it is not an equation (no = sign). we are just simplifying it. so -1*61 + -1*-9y is the same as -61 + 9y do you understand?

OpenStudy (anonymous):

yes thanks

OpenStudy (phi):

Ok, now back to the equation -1*(61-9y) + y = -1 can you distribute the -1 ? Don't do too much. One step at a time.

OpenStudy (anonymous):

-52+y=-1

OpenStudy (anonymous):

-1(-52)+y=-1

OpenStudy (phi):

"distribute the -1" That means multiply every term inside the parentheses by -1. -1*(61-9y) = -1*61 + -1*-9y try again